OpenStudy (anonymous):

Find the equation of the line that passes through the points (4, 8) and (8, 11).

OpenStudy (ash2326):

Bailey first find slope and then satisfy any of the two points to find the equation of the line. \[slope (m)=\frac{y2-y1}{x2-x1}\] and \[y=mx +c\] find m and then find c

OpenStudy (ash2326):

Can you do it ?

OpenStudy (anonymous):

nvm :) i got it its y=3/4x-5

OpenStudy (ash2326):

Bailey it'll be \[y=\frac{3}{4} x+c\] let's use 4,8 here \[8=\frac{3}{4}\times 4+c\] we get \[8=3+c\] so \[c=5\] so \[y=\frac{3}{4}x+5\]